\begin{thebibliography}{49}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Ahmed and Wahed(2020)]{ahmed2020democratization}
N.~Ahmed and M.~Wahed.
\newblock The de-democratization of ai: Deep learning and the compute divide in
  artificial intelligence research.
\newblock \emph{arXiv preprint arXiv:2010.15581}, 2020.

\bibitem[Bach and Jordan(2002)]{bach2002kernel}
F.~R. Bach and M.~I. Jordan.
\newblock Kernel independent component analysis.
\newblock \emph{Journal of Machine Learning Research}, 3\penalty0
  (Jul):\penalty0 1--48, 2002.

\bibitem[Chen et~al.(2019)Chen, Medini, Farwell, Gobriel, Tai, and
  Shrivastava]{chen2019slide}
B.~Chen, T.~Medini, J.~Farwell, S.~Gobriel, C.~Tai, and A.~Shrivastava.
\newblock Slide: In defense of smart algorithms over hardware acceleration for
  large-scale deep learning systems.
\newblock \emph{arXiv preprint arXiv:1903.03129}, 2019.

\bibitem[Chicco and Jurman(2020)]{chicco2020machine}
D.~Chicco and G.~Jurman.
\newblock Machine learning can predict survival of patients with heart failure
  from serum creatinine and ejection fraction alone.
\newblock \emph{BMC Medical Informatics and Decision Making}, 20\penalty0
  (1):\penalty0 16, 2020.

\bibitem[Cios et~al.(2001)Cios, Kurgan, and Goodenday]{Cios:2001}
K.~J. Cios, L.~A. Kurgan, and L.~S. Goodenday.
\newblock Spectf heart data set - {UCI} machine learning repository, 2001.
\newblock URL \url{https://archive.ics.uci.edu/ml/datasets/SPECTF+Heart}.

\bibitem[Garvey(2018)]{garvey2018framework}
C.~Garvey.
\newblock A framework for evaluating barriers to the democratization of
  artificial intelligence.
\newblock In \emph{Thirty-Second AAAI Conference on Artificial Intelligence},
  2018.

\bibitem[Gaur et~al.(2021)Gaur, Faldu, and Sheth]{gaur2021semantics}
M.~Gaur, K.~Faldu, and A.~Sheth.
\newblock Semantics of the black-box: Can knowledge graphs help make deep
  learning systems more interpretable and explainable?
\newblock \emph{IEEE Internet Computing}, 25\penalty0 (1):\penalty0 51--59,
  2021.

\bibitem[He et~al.(2016)He, Zhang, Ren, and Sun]{he2016deep}
K.~He, X.~Zhang, S.~Ren, and J.~Sun.
\newblock Deep residual learning for image recognition.
\newblock In \emph{Proceedings of the IEEE Conference on Computer Vision and
  Pattern Recognition}, pages 770--778, 2016.

\bibitem[Hsu et~al.(2015)Hsu, Mullen, Jung, and Cauwenberghs]{hsu2015real}
S.-H. Hsu, T.~R. Mullen, T.-P. Jung, and G.~Cauwenberghs.
\newblock Real-time adaptive eeg source separation using online recursive
  independent component analysis.
\newblock \emph{IEEE Transactions on Neural Systems and Rehabilitation
  Engineering}, 24\penalty0 (3):\penalty0 309--319, 2015.

\bibitem[Hsu et~al.(2018)Hsu, Pion-Tonachini, Palmer, Miyakoshi, Makeig, and
  Jung]{hsu2018modeling}
S.-H. Hsu, L.~Pion-Tonachini, J.~Palmer, M.~Miyakoshi, S.~Makeig, and T.-P.
  Jung.
\newblock Modeling brain dynamic state changes with adaptive mixture
  independent component analysis.
\newblock \emph{NeuroImage}, 183:\penalty0 47--61, 2018.

\bibitem[Hyv{\"a}rinen and Oja(2000)]{hyvarinen2000independent}
A.~Hyv{\"a}rinen and E.~Oja.
\newblock Independent component analysis: algorithms and applications.
\newblock \emph{Neural Networks}, 13\penalty0 (4-5):\penalty0 411--430, 2000.

\bibitem[Jacot et~al.(2018)Jacot, Gabriel, and Hongler]{jacot2018neural}
A.~Jacot, F.~Gabriel, and C.~Hongler.
\newblock Neural tangent kernel: Convergence and generalization in neural
  networks.
\newblock In \emph{Advances in Neural Information Processing Systems}, pages
  8571--8580, 2018.

\bibitem[James and Hesse(2004)]{james2004independent}
C.~J. James and C.~W. Hesse.
\newblock Independent component analysis for biomedical signals.
\newblock \emph{Physiological Measurement}, 26\penalty0 (1):\penalty0 R15,
  2004.

\bibitem[Javidi et~al.(2011)Javidi, Took, and Mandic]{javidi2011fast}
S.~Javidi, C.~C. Took, and D.~P. Mandic.
\newblock Fast independent component analysis algorithm for quaternion valued
  signals.
\newblock \emph{IEEE Transactions on Neural Networks}, 22\penalty0
  (12):\penalty0 1967--1978, 2011.

\bibitem[Kobayashi et~al.(2019)Kobayashi, Ishibashi, and
  Kobayashi]{kobayashi2019will}
Y.~Kobayashi, M.~Ishibashi, and H.~Kobayashi.
\newblock How will “democratization of artificial intelligence” change the
  future of radiologists?
\newblock \emph{Japanese Journal of Radiology}, 37\penalty0 (1):\penalty0
  9--14, 2019.

\bibitem[Krizhevsky et~al.(2012)Krizhevsky, Sutskever, and
  Hinton]{krizhevsky2012imagenet}
A.~Krizhevsky, I.~Sutskever, and G.~E. Hinton.
\newblock Imagenet classification with deep convolutional neural networks.
\newblock \emph{Advances in Neural Information Processing Systems},
  25:\penalty0 1097--1105, 2012.

\bibitem[LeCun et~al.(1995)LeCun, Bengio, et~al.]{lecun1995convolutional}
Y.~LeCun, Y.~Bengio, et~al.
\newblock Convolutional networks for images, speech, and time series.
\newblock \emph{The Handbook of Brain Theory and Neural Networks},
  3361\penalty0 (10):\penalty0 1995, 1995.

\bibitem[Li and Adali(2010)]{li2010complex}
X.-L. Li and T.~Adali.
\newblock Complex independent component analysis by entropy bound minimization.
\newblock \emph{IEEE Transactions on Circuits and Systems I: Regular Papers},
  57\penalty0 (7):\penalty0 1417--1430, 2010.

\bibitem[Lim(1999)]{Lim:1999}
T.-S. Lim.
\newblock Haberman's survival data set - {UCI} machine learning repository,
  1999.
\newblock URL
  \url{https://archive.ics.uci.edu/ml/datasets/Haberman's+Survival}.

\bibitem[Little et~al.(2007)Little, McSharry, Roberts, Costello, and
  Moroz]{little2007exploiting}
M.~A. Little, P.~E. McSharry, S.~J. Roberts, D.~A.~E. Costello, and I.~M.
  Moroz.
\newblock Exploiting nonlinear recurrence and fractal scaling properties for
  voice disorder detection.
\newblock \emph{Biomedical Engineering Online}, 6\penalty0 (1):\penalty0 23,
  2007.

\bibitem[Liu et~al.(2021)Liu, Mocanu, Matavalam, Pei, and
  Pechenizkiy]{liu2021sparse}
S.~Liu, D.~C. Mocanu, A.~R.~R. Matavalam, Y.~Pei, and M.~Pechenizkiy.
\newblock Sparse evolutionary deep learning with over one million artificial
  neurons on commodity hardware.
\newblock \emph{Neural Computing and Applications}, 33\penalty0 (7):\penalty0
  2589--2604, 2021.

\bibitem[McKeown et~al.(2003)McKeown, Hansen, and
  Sejnowsk]{mckeown2003independent}
M.~J. McKeown, L.~K. Hansen, and T.~J. Sejnowsk.
\newblock Independent component analysis of functional mri: what is signal and
  what is noise?
\newblock \emph{Current Opinion in Neurobiology}, 13\penalty0 (5):\penalty0
  620--629, 2003.

\bibitem[Nguyen et~al.(2018)Nguyen, Kalra, Jun, and Kim]{nguyen2018novel}
H.~M. Nguyen, G.~Kalra, T.~J. Jun, and D.~Kim.
\newblock A novel echo state network model using bayesian ridge regression and
  independent component analysis.
\newblock In \emph{International Conference on Artificial Neural Networks},
  pages 24--34. Springer, 2018.

\bibitem[Parisi(2014{\natexlab{a}})]{parisi2014exploiting}
L.~Parisi.
\newblock Exploiting kinetic and kinematic data to plot cyclograms for managing
  the rehabilitation process of bkas by applying neural networks.
\newblock \emph{International Journal of Medical, Health, Pharmaceutical and
  Biomedical Engineering}, 2014{\natexlab{a}}.

\bibitem[Parisi(2014{\natexlab{b}})]{parisi2014neural}
L.~Parisi.
\newblock Neural networks for distinguishing the performance of two hip joint
  implants on the basis of hip implant side and ground reaction force.
\newblock \emph{International Journal of Medical, Health, Pharmaceutical and
  Biomedical Engineering}, 2014{\natexlab{b}}.

\bibitem[Parisi(2019)]{parisi2019machine}
L.~Parisi.
\newblock \emph{Machine Learning-based Feature Selection and Optimisation for
  Clinical Decision Support Systems. Optimal Data-driven Feature Selection
  Methods for Binary and Multi-class Classification Problems: Towards a Minimum
  Viable Solution for Predicting Early Diagnosis and Prognosis}.
\newblock PhD thesis, University of Bradford, 2019.

\bibitem[Parisi(2020)]{parisi2020m}
L.~Parisi.
\newblock m-arcsinh: An efficient and reliable function for svm and mlp in
  scikit-learn.
\newblock \emph{arXiv preprint arXiv:2009.07530}, 2020.

\bibitem[Parisi and Manaog(2016)]{parisi2016preliminary}
L.~Parisi and M.~L. Manaog.
\newblock Preliminary validation of the lagrangian support vector machine
  learning classifier as clinical decision-making support tool to aid
  prediction of prognosis in patients with hepatitis.
\newblock In \emph{The 16th International Conference on Biomedical Engineering,
  National University of Singapore (NUS)}, 2016.

\bibitem[Parisi and Manaog(2017{\natexlab{a}})]{parisi2017importance}
L.~Parisi and M.~L. Manaog.
\newblock The importance of selecting appropriate k-fold cross-validation and
  training algorithms in improving postoperative discharge decision-making via
  artificial intelligence.
\newblock In \emph{2017 AUT mathematical Sciences Symposium}, volume~1,
  page~16, 2017{\natexlab{a}}.

\bibitem[Parisi and Manaog(2017{\natexlab{b}})]{parisi2017minimum}
L.~Parisi and M.~L. Manaog.
\newblock A minimum viable machine learning-based speech processing solution
  for facilitating early diagnosis of parkinson’s disease.
\newblock In \emph{MATLAB Conference 2017, Auckland, New Zealand},
  2017{\natexlab{b}}.

\bibitem[Parisi and RaviChandran(2020)]{parisi2020evolutionary}
L.~Parisi and N.~RaviChandran.
\newblock Evolutionary feature transformation to improve prognostic prediction
  of hepatitis.
\newblock \emph{Knowledge-Based Systems}, 200:\penalty0 106012, 2020.

\bibitem[Parisi et~al.(2015)Parisi, Biggs, Whatling, and Holt]{parisi2015novel}
L.~Parisi, P.~R. Biggs, G.~M. Whatling, and C.~A. Holt.
\newblock A novel comparison of artificial intelligence methods for diagnosing
  knee osteoarthritis.
\newblock In \emph{XXV Congress of the International Society of Biomechanics},
  2015.

\bibitem[Parisi et~al.(2018{\natexlab{a}})Parisi, RaviChandran, and
  Manaog]{parisi2018decision}
L.~Parisi, N.~RaviChandran, and M.~L. Manaog.
\newblock Decision support system to improve postoperative discharge: A novel
  multi-class classification approach.
\newblock \emph{Knowledge-Based Systems}, 152:\penalty0 1--10,
  2018{\natexlab{a}}.

\bibitem[Parisi et~al.(2018{\natexlab{b}})Parisi, RaviChandran, and
  Manaog]{parisi2018feature}
L.~Parisi, N.~RaviChandran, and M.~L. Manaog.
\newblock Feature-driven machine learning to improve early diagnosis of
  parkinson's disease.
\newblock \emph{Expert Systems with Applications}, 110:\penalty0 182--190,
  2018{\natexlab{b}}.

\bibitem[Parisi et~al.(2020{\natexlab{a}})Parisi, Neagu, Ma, and
  Campean]{parisi2020qrelu}
L.~Parisi, D.~Neagu, R.~Ma, and F.~Campean.
\newblock Qrelu and m-qrelu: Two novel quantum activation functions to aid
  medical diagnostics.
\newblock \emph{arXiv preprint arXiv:2010.08031}, 2020{\natexlab{a}}.

\bibitem[Parisi et~al.(2020{\natexlab{b}})Parisi, RaviChandran, and
  Manaog]{parisi2020novel}
L.~Parisi, N.~RaviChandran, and M.~L. Manaog.
\newblock A novel hybrid algorithm for aiding prediction of prognosis in
  patients with hepatitis.
\newblock \emph{Neural Computing and Applications}, 32\penalty0 (8):\penalty0
  3839--3852, 2020{\natexlab{b}}.

\bibitem[Parisi et~al.(2021{\natexlab{a}})Parisi, Ma, RaviChandran, and
  Lanzillotta]{parisi2021hyper}
L.~Parisi, R.~Ma, N.~RaviChandran, and M.~Lanzillotta.
\newblock hyper-sinh: An accurate and reliable function from shallow to deep
  learning in tensorflow and keras.
\newblock \emph{Machine Learning with Applications}, page 100112,
  2021{\natexlab{a}}.

\bibitem[Parisi et~al.(2021{\natexlab{b}})Parisi, Ma, RaviChandran, and
  Lanzillotta]{parisimachine}
L.~Parisi, R.~Ma, N.~RaviChandran, and M.~Lanzillotta.
\newblock Machine learning with applications.
\newblock \emph{Machine Learning with Applications}, 2021{\natexlab{b}}.

\bibitem[Parisi et~al.(2021{\natexlab{c}})Parisi, Ma, Zaernia, and
  Youseffi]{parisietayper}
L.~Parisi, R.~Ma, A.~Zaernia, and M.~Youseffi.
\newblock Hyper-sinh-convolutional neural network for early detection of
  parkinson’s disease from spiral drawings.
\newblock \emph{WSEAS Transactions on Computer Research}, 2021{\natexlab{c}}.

\bibitem[Parisi et~al.(2021{\natexlab{d}})Parisi, Ma, Zaernia, and
  Youseffi]{parisim}
L.~Parisi, R.~Ma, A.~Zaernia, and M.~Youseffi.
\newblock m-ark-support vector machine for early detection of parkinson’s
  disease from speech signals.
\newblock \emph{International Journal of Mathematics and Computers in
  Simulation}, 2021{\natexlab{d}}.

\bibitem[Pedregosa et~al.(2011)Pedregosa, Varoquaux, Gramfort, Michel, Thirion,
  Grisel, Blondel, Prettenhofer, Weiss, Dubourg, Vanderplas, Passos,
  Cournapeau, Brucher, Perrot, and Duchesnay]{scikit-learn}
F.~Pedregosa, G.~Varoquaux, A.~Gramfort, V.~Michel, B.~Thirion, O.~Grisel,
  M.~Blondel, P.~Prettenhofer, R.~Weiss, V.~Dubourg, J.~Vanderplas, A.~Passos,
  D.~Cournapeau, M.~Brucher, M.~Perrot, and E.~Duchesnay.
\newblock Scikit-learn: Machine learning in {P}ython.
\newblock \emph{Journal of Machine Learning Research}, 12:\penalty0 2825--2830,
  2011.

\bibitem[Rasmussen et~al.(2020)Rasmussen, Siersma, Malmqvist, and
  Brodersen]{rasmussen2020psychosocial}
J.~F. Rasmussen, V.~Siersma, J.~Malmqvist, and J.~Brodersen.
\newblock Psychosocial consequences of false positives in the danish lung
  cancer ct screening trial: a nested matched cohort study.
\newblock \emph{BMJ Open}, 10\penalty0 (6):\penalty0 e034682, 2020.

\bibitem[Rumelhart et~al.(1986)Rumelhart, Hinton, and
  Williams]{rumelhart1986learning}
D.~E. Rumelhart, G.~E. Hinton, and R.~J. Williams.
\newblock Learning representations by back-propagating errors.
\newblock \emph{Nature}, 323\penalty0 (6088):\penalty0 533--536, 1986.

\bibitem[Samek and M{\"u}ller(2019)]{samek2019towards}
W.~Samek and K.-R. M{\"u}ller.
\newblock Towards explainable artificial intelligence.
\newblock In \emph{Explainable AI: Interpreting, explaining and visualizing
  deep learning}, pages 5--22. Springer, 2019.

\bibitem[Shen et~al.(2009)Shen, Jegelka, and Gretton]{shen2009fast}
H.~Shen, S.~Jegelka, and A.~Gretton.
\newblock Fast kernel-based independent component analysis.
\newblock \emph{IEEE Transactions on Signal Processing}, 57\penalty0
  (9):\penalty0 3498--3511, 2009.

\bibitem[Simonyan and Zisserman(2014)]{simonyan2014very}
K.~Simonyan and A.~Zisserman.
\newblock Very deep convolutional networks for large-scale image recognition.
\newblock \emph{arXiv preprint arXiv:1409.1556}, 2014.

\bibitem[Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, G.,
  Kaiser, and Polosukhin]{vaswani2017attention}
A.~Vaswani, N.~Shazeer, N.~Parmar, J.~Uszkoreit, L.~Jones, Aidan~N. G.,
  {\L}.~Kaiser, and I.~Polosukhin.
\newblock Attention is all you need.
\newblock In \emph{Advances in Neural Information Processing Systems}, pages
  5998--6008, 2017.

\bibitem[Vert and Vert(2006)]{vert2006consistency}
R.~Vert and J.-P. Vert.
\newblock Consistency and convergence rates of one-class svms and related
  algorithms.
\newblock \emph{Journal of Machine Learning Research}, 7\penalty0
  (May):\penalty0 817--854, 2006.

\bibitem[Wolberg et~al.(1995)Wolberg, Street, and Mangasarian]{Wolberg:1995}
W.~H. Wolberg, W.~N. Street, and O.~L. Mangasarian.
\newblock Breast cancer wisconsin (diagnostic) data set - {UCI} machine
  learning repository, 1995.
\newblock URL
  \url{https://archive.ics.uci.edu/ml/datasets/Breast+Cancer+Wisconsin+(Diagnostic)}.

\end{thebibliography}

\documentclass[twoside,11pt]{article}

\usepackage{jmlr2e}
\usepackage{graphicx}
\usepackage{float}
\usepackage{listings}
\usepackage{mathtools}
\usepackage{pgfplots}
\usepackage[super]{nth}

\pgfkeys{/pgf/declare function={arcsinh(\x) = ln(\x + sqrt(\x^2+1));}}
\pgfplotsset{compat=1.14}

\ShortHeadings{m-ar-K-FastICA: Reliable Feature Extraction in scikit-learn}{Luca Parisi, \emph{PhD, MBA Candidate}}
\firstpageno{1}

\begin{document}

\title{m-ar-K-Fast Independent Component Analysis}

\author{\name Luca Parisi \email luca.parisi@ieee.org \\
       \addr Coventry, United Kingdom \\
       PhD in Machine Learning for Clinical Decision Support Systems \\ 
       MBA Candidate with Artificial Intelligence Specialism
       }

\editor{}

\maketitle


\begin{abstract}
This study presents the m-arcsinh Kernel ('m-ar-K') Fast Independent Component Analysis ('FastICA') method ('m-ar-K-FastICA') for feature extraction. The kernel trick has enabled dimensionality reduction techniques to capture a higher extent of non-linearity in the data; however, reproducible, open-source kernels to aid with feature extraction are still limited and may not be reliable when projecting features from entropic data. The m-ar-K function, freely available in Python and compatible with its open-source library 'scikit-learn', is hereby coupled with FastICA to achieve more reliable feature extraction in presence of a high extent of randomness in the data, reducing the need for pre-whitening. Different classification tasks were considered, as related to five (N = 5) open access datasets of various degrees of information entropy, available from scikit-learn and the University California Irvine (UCI) Machine Learning repository. Experimental results demonstrate improvements in the classification performance brought by the proposed feature extraction. The novel m-ar-K-FastICA dimensionality reduction approach is compared to the 'FastICA' gold standard method, supporting its higher reliability and computational efficiency, regardless of the underlying uncertainty in the data.

\end{abstract}

\vskip 0.1in

\begin{keywords}
  m-arcsinh, Kernel, m-ar-K-FastICA, Independent Component Analysis, ICA, FastICA, scikit-learn
\end{keywords}

\newpage


\vskip 0.3in

\section{Introduction}

\vskip 0.15in

Despite theoretical advances in the Independent Component Analysis (ICA) feature extraction method, suitable for dimensionality reduction in presence of non-Gaussian/non-normal data distributions, such as the FastICA~\citep{hyvarinen2000independent}~\citep{javidi2011fast}, the Kernel ICA (KICA)~\citep{bach2002kernel}, the Fast Kernel-based ICA (FastKICA)~\citep{shen2009fast}, and the Adaptive Mixture ICA (AMICA)~\citep{hsu2018modeling}, usable, explainable, reproducible and replicable variations of the ICA method that improve feature extraction are still limited. In fact, the open-source Python library named 'scikit-learn'~\citep{scikit-learn} leverages the FastICA algorithm as the only ICA-related variation available, which uses the ~\emph{logcosh} as the G function to approximate to neg-entropy~\citep{nguyen2018novel}.
\\

However, such variants of ICA, even with appropriate data pre-whitening~\citep{james2004independent}~\citep{hsu2015real}, may not be reliable when performing dimensionality reduction of entropic data~\citep{mckeown2003independent}~\citep{li2010complex}, characterised by a high extent of randomness or uncertainty. When applied prior to classification, they can lead to slow or lack of convergence~\citep{vert2006consistency}~\citep{jacot2018neural}, due to trapping at local \emph{minima}~\citep{parisi2014exploiting}~\citep{parisi2016preliminary}~\citep{parisi2017importance}~\citep{parisi2017minimum}. Furthermore, despite recent advancements in kernel and activation functions across both shallow and deep learning~\citep{parisi2020qrelu}~\citep{parisietayper}~\citep{parisimachine}~\citep{parisi2021hyper}, it is still challenging to derive a kernel function that can aid feature extraction methods to generalise in presence of non-Gaussian and entropic data~\citep{parisi2018decision}~\citep{parisi2020novel}~\citep{parisi2020evolutionary}.
\\

To address these challenges, over the past two decades, there has been an increasing trend in the Artificial Intelligence (AI) community to generate deep models that incorporate feature extraction within their architectures, starting from the Convolutional Neural Network (CNN)~\citep{lecun1995convolutional} to deep CNN-like variations~\citep{krizhevsky2012imagenet}~\citep{simonyan2014very}~\citep{he2016deep}, and, more recently, even deeper models, such as Transformer-type of neural networks ~\citep{vaswani2017attention}. Nevertheless, this school of thought attempts to handle entropy in the data intrinsically, by training deep models that are capable of automated feature extraction on expensive hardware, e.g., Graphics and Tensor Processing Units. 
\\

This trendy reliance on deep models and costly hardware has transformed AI into an elitist field, albeit with a very few exceptions~\citep{chen2019slide}~\citep{liu2021sparse}, as it demands expensive computing resources for predictive analytics. Furthermore, it has jeopardised the explainability of classifications~\citep{gaur2021semantics}, as features are extracted and transformed within deep model architectures, which, \emph{per se}, are already not straightforward to describe to end users. In fact, these limitations have hindered a wide and transparent application of deep learning models in regulated industries, such as finance and healthcare~\citep{samek2019towards}.
\\

To contribute to the democratisation of AI~\citep{garvey2018framework}~\citep{kobayashi2019will}~\citep{ahmed2020democratization}, thus enabling resource-constrained environments to leverage extrinsic and explainable feature extraction methods for AI to have a wider positive impact in society, in this study, a novel kernel-based dimensionality reduction method is proposed to reduce the uncertainty in the data conveyed to improve classification performance. Written in Python and made freely available in 'scikit-learn'~\citep{scikit-learn} for the 'FastICA' class, the proposed feature extraction method is presented as a more reliable approach than the gold standard FastICA method. By virtue of its liberal license, this kernel-based dimensionality reduction method is broadly disseminated within the free software Python library 'scikit-learn'~\citep{scikit-learn}, and it can be used for both academic research and commercial purposes.\\

\newpage


\vskip 0.3in

\section{Methods}

\vskip 0.15in

\subsection{Datasets leveraged from scikit-learn and the UCI ML repository}

\vskip 0.15in

In this study, the following datasets from \emph{The University California Irvine (UCI) ML repository} were used:

\begin{itemize}

\item \href{https://archive.ics.uci.edu/ml/machine-learning-databases/parkinsons/parkinsons.data}{‘Parkinson’s’ dataset}~\citep{little2007exploiting}, which has 23 features corresponding to 195 biomedical voice measurements from 31 people, 23 with Parkinson's disease (PD), to help in detecting PD from speech signals.

\item \href{https://archive.ics.uci.edu/ml/machine-learning-databases/haberman/haberman.data}{‘Haberman survival’ dataset}~\citep{Lim:1999}, with three features (age of patient at time of operation; patient's year of operation; number of positive axillary nodes detected) to predict whether 306 patients who had undergone surgery for breast cancer would have died within 5 years of follow up or survived for longer.

\item \href{https://archive.ics.uci.edu/ml/machine-learning-databases/00519/heart_failure_clinical_records_dataset.csv}{‘Heart failure clinical records’ dataset}~\citep{chicco2020machine}, to predict whether a patient was deceased during the follow-up period, based on 13 clinical features from medical records of 299 patients who had heart failure.

\item ‘SPECTF’ dataset~\citep{Cios:2001}, which has 267 images collected via a cardiac Single Proton Emission Computed Tomography (SPECT), describing whether each patient has a physiological or pathophysiological heart based on 44 features, as per the following two data splits or partitions:

\begin{itemize}
\item \href{https://archive.ics.uci.edu/ml/machine-learning-databases/spect/SPECTF.train}{training data partition} ('SPECTF.train' file), with 80 images.
\item \href{https://archive.ics.uci.edu/ml/machine-learning-databases/spect/SPECTF.test}{testing data partition} ('SPECTF.test' file), which has 187 images.
\end{itemize}

\end{itemize}

Moreover, the following dataset from \emph{scikit-learn} was leveraged for a further evaluation:

\begin{itemize}

\item 'Breast cancer Wisconsin (diagnostic)' dataset~\citep{Wolberg:1995}, having 30 characteristics of cell \emph{nuclei} from 569 digitised images of a fine needle aspirate of breast masses, to detect whether they correspond to either malignant or benign breast cancer.

\end{itemize}

\vskip 0.15in

\subsection{Baseline MLP model, activation functions, and hyperparameters}

\vskip 0.15in

The purpose of this study is not to devise the most optimised, best-performing classifier for any of the classification tasks involved in section 2.1, but to develop a novel reliable kernel-based feature extraction method. Thus, whilst applying early stopping to avoid overfitting, a baseline Multi-Layer Perceptron (MLP)~\citep{rumelhart1986learning}~\citep{parisi2014neural}~\citep{parisi2015novel}~\citep{parisi2018feature}~\citep{parisi2019machine} model was leveraged using various benchmark activation functions (m-arcsinh~\citep{parisi2020m}~\citep{parisim}, identity, tanh~\citep{jacot2018neural}, and ReLU) and with the following hyperparameters for all classification tasks in section 2.1:

\begin{itemize}

\item MLP-related hyperparameters:
\begin{itemize}
\item \emph{'random\_state'} = 1.
\item \emph{'max\_iter'} = 250, where 'max\_iter' is the maximum number of iterations.
\end{itemize}

\end{itemize}

Listing 1 provides the snippet of code in Python to use an MLP with different activation functions available in 'scikit-learn'~\citep{scikit-learn}, including the 'm-arcsinh', following feature extraction.

\lstset{language=Python}
\lstset{frame=lines}
\lstset{caption={MLP with different activation functions available in 'scikit-learn'~\citep{scikit-learn}, including the 'm-arcsinh'.}}
\lstset{label={lst:code_direct_1}}
\lstset{basicstyle=\footnotesize}
\begin{lstlisting}

from sklearn.neural_network import MLPClassifier

for activation in ('m-arcsinh', 'identity', 'tanh', 'relu'):
    classifier =  MLPClassifier(activation=activation, 
    random_state=1, max_iter=1000)

\end{lstlisting}

Except for the ‘SPECTF’ dataset~\citep{Cios:2001}, which is already provided in two separate partitions for training and testing (see section 2.1), the other four datasets were split into 80\% for training and 20\% for testing via \emph{'train\_test\_split'} in 'scikit-learn'~\citep{scikit-learn} from \emph{'sklearn.model\_selection'}, without randomisation (\emph{'shuffle'=False}) to enable reproducibility of the results presented in this study.

\vskip 0.15in

\subsection{m-arK-FastICA: A novel reliable feature extraction approach}

Coupling FastICA with the modified ('m-') inverse hyperbolic sine function ('arcsinh') or 'm-arcsinh' function that was found to generalise as a kernel and activation for optimal separating hyperplane- and neural network-based classifiers~\citep{parisi2020m}~\citep{parisim}, the proposed feature extraction method seeks to 1) handle the non-linearity in the data further via an improved transformation in the G function that is leveraged in the approximation to neg-entropy, 2) cope with entropic data by coupling the m-arcsinh Kernel ('m-ar-K') with FastICA, thus achieving the hybrid method 'm-ar-K-FastICA', and 3) improve classification of non-linearly separable and entropic input data into target classes by fulfilling points 1) and 2).

Via a weighted interaction effect between the arcsinh and the slightly non-linear characteristic of the squared root function, the m-ar-K~\citep{parisi2020m}~\citep{parisim} can be represented as follows:

\vspace{1em}
\begin{math}
y = arcsinh(x) \times \frac{1}{3} \times \frac{1}{4} \times \sqrt{\left|x\right|} = arcsinh(x) \times \frac{1}{12} \times  \sqrt{\left|x\right|}\hspace{13.5em}(1)
\end{math}
\vspace{2em}

\begin{tikzpicture}
\begin{axis}[ymin=-0.1, ymax=0.1, enlargelimits=false]
\addplot [domain=-1:1, samples=101,unbounded coords=jump]{(arcsinh(x)/3)*(sqrt(abs(x))/4)};
\end{axis}
\end{tikzpicture}

\vspace{2em}

The derivative of m-ar-K~\citep{parisi2020m}~\citep{parisim} is described as follows:

\vspace{1em}
\begin{math}
{\frac{dy}{dx} = \sqrt{\left|x\right|}} \times \frac{1}{12\times\sqrt{x^2+1}}  +\  \frac{x\ \times arcsinh(x)}{24\times\left|x\right|^\frac{3}{2}\ }\hspace{21.5em}(2)
\end{math}
\hspace{36.5em}(2)
\vspace{2em}

\begin{tikzpicture}
\begin{axis}[ymin=0, ymax=0.1, enlargelimits=false]
\addplot [domain=-1:1, samples=101,unbounded coords=jump]{sqrt(abs(x))/(12*sqrt(x^2+1)) + (x*arcsinh(x))/(24*abs(x)^(3/2)))};
\end{axis}
\end{tikzpicture}

\vspace{2em}

To enable reproducibility of the results obtained in this study, considering that the number of components (~\emph{n\_components}) were set based on the dimensionality and the level of entropy of the data, the following parameters were used in both the FastICA and the proposed m-ar-K-FastICA method to extract features from:

\begin{itemize}
\item All datasets:
\begin{itemize}
\item \emph{'random\_state'} = 42.
\end{itemize}
\end{itemize}

\begin{itemize}
\item The ‘Haberman survival’~\citep{Lim:1999} and the ‘Heart failure clinical records’ datasets~\citep{chicco2020machine}:
\begin{itemize}
\item \emph{'n\_components'} = 2.
\end{itemize}
\end{itemize}

\begin{itemize}
\item The ‘Parkinson's’ dataset~\citep{little2007exploiting}:
\begin{itemize}
\item \emph{'n\_components'} = 8.
\end{itemize}
\end{itemize}

\begin{itemize}
\item The ‘Breast cancer Wisconsin (diagnostic)’ dataset~\citep{Wolberg:1995}:
\begin{itemize}
\item \emph{'n\_components'} = 16.
\end{itemize}
\end{itemize}

\begin{itemize}
\item The ‘SPECTF’ dataset~\citep{Cios:2001}:
\begin{itemize}
\item \emph{'n\_components'} = 43.
\end{itemize}
\end{itemize}

Listing 2 provides the snippet of code in Python that implements the proposed m-arcsinh Kernel (m-ar-K) (1) and its derivative (2) coupled in the 'FastICA' class as a G function leveraged to approximate to neg-entropy in 'scikit-learn'~\citep{scikit-learn}.

\lstset{language=Python}
\lstset{frame=lines}
\lstset{caption={Coupling the m-arcsinh kernel ('m-ar-K') and its derivative as a G function in the approximation to neg-entropy in 'scikit-learn'~\citep{scikit-learn}.}}
\lstset{label={lst:code_direct_4}}
\lstset{basicstyle=\footnotesize}
\begin{lstlisting}

import numpy as np


def _m_arcsinh(x, fun_args):
    """Compute the m-arcsinh kernel function and its derivative 
    in place.
    
    It exploits the fact that the derivative is a simple function of the 
    output value from the m-arcsinh kernel.

    Parameters
    ----------
    x: {array-like, sparse matrix}, shape (n_samples, n_features)
        The input data.

    Returns
    -------
    A tuple containing the value of the function, and that of its derivative.
    """

    return (1/3*np.arcsinh(x))*(1/4*np.sqrt(np.abs(x))),
           (np.sqrt(np.abs(x))/(12*np.sqrt(x**2+1)) 
           + (x*np.arcsinh(x))/(24*np.abs(x)**(3/2))).mean(axis=-1)




# Using the m-arcsinh kernel as a G function ('fun') inside the class 'FastICA'
# for feature extraction (using 16 components and fixed random state set to 42 
# for reproducibility in the example below)

if self.fun == 'logcosh':
    g = _logcosh
elif self.fun == 'exp':
    g = _exp
elif self.fun == 'cube':
    g = _cube
elif self.fun == 'm_arcsinh':
    g = _m_arcsinh
elif callable(self.fun):
    def g(x, fun_args):
        return self.fun(x, **fun_args)
else:
    exc = ValueError if isinstance(self.fun, str) else TypeError
    raise exc(
        "Unknown function %r;"
        " should be one of 'logcosh', 'exp', 'cube', 'm_arcsinh', or callable"
        % self.fun
    )


from sklearn.decomposition import FastICA
transformer = FastICA(n_components=16, random_state=42, fun='m_arcsinh')
transformed_inputs = transformer.fit_transform(inputs)

\end{lstlisting}

\vskip 0.15in

\subsection{Performance evaluation}

\vskip 0.15in

Further to feature extraction as per section 2.3, the accuracy of the MLP classifier in section 2.2 using different activation functions (m-arcsinh~\citep{parisi2020m}~\citep{parisim}, identity, tanh~\citep{jacot2018neural}, and ReLU) on the datasets described in section 2.1, was evaluated via the \emph{'accuracy\_score'} available in 'scikit-learn'~\citep{scikit-learn} from \emph{'sklearn.metrics'}. 
The reliability of such classifiers was assessed via the weighted average of the precision, recall and F1-score computed via the \emph{'classification\_report'}, also available in 'scikit-learn'~\citep{scikit-learn} from \emph{'sklearn.metrics'}.
\\
To understand what classification accuracy and reliability are, and how they can be evaluated, please refer to the following studies:~\citep{parisi2018decision},~\citep{parisi2018feature},~\citep{parisi2020novel},~\citep{parisi2020evolutionary}.

Moreover, the computational cost of the classifier, to quantify the impact of using the proposed feature extraction method (m-ar-K-FastICA) as opposed to the current gold standard one (FastICA), was assessed via the training time in seconds. Experiments were run on an AMD E2-9000 Radeon R2 processor, 1.8 GHz and 4 GB DDR4 RAM.

\newpage


\section{Results}

\vskip 0.15in

Experimental results demonstrate the overall higher reliability achieved via the proposed m-ar-K-FastICA kernel-based dimensionality reduction method for MLP-based classification, including a shorter training time, as follows:
\begin{itemize}
\item Higher reliability on 2 of 5 datasets evaluated (Tables 1 and 3).
\item The same classification performance on 3 of 5 datasets assessed (Tables 2, 4, and 5).
\item Shorter training time for 3 of 5 datasets evaluated (Tables 2-4).
\end{itemize}

\vskip 0.145in


\textbf{Table 1} Results on performance evaluation of baseline (non-optimised) Multi-Layer Perceptron (MLP) in scikit-learn with different kernel functions, to compare the impact of using either the gold standard FastICA feature extraction algorithm or the proposed m-ar-K-FastICA method. The performance of this classifier was assessed on the ‘Breast cancer Wisconsin (diagnostic)’ dataset~\citep{Wolberg:1995} available in scikit-learn.

\begin{table}[H]
\resizebox{\textwidth}{!}{%
\begin{tabular}{llllllll}
\textbf{Classifier} &
  \textbf{Kernel} &
  \textbf{Feature extraction} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Training time \\ \\ (s)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Accuracy \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Precision \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Recall \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}F1-score \\ \\ (0-1)\end{tabular}} \\
MLP & m-arcsinh & FastICA & 4.71 & 0.95 & 0.95 & 0.95 & 0.95 \\
MLP & Identity & FastICA & 1.64 & 0.95 & 0.95 & 0.95 & 0.95 \\
MLP & tanh & FastICA & 2.97 & 0.97 & 0.97 & 0.97 & 0.97 \\
MLP & ReLU & FastICA & 1.87 & 0.95 & 0.95 & 0.95 & 0.95 \\
MLP & m-arcsinh & \textbf{m-ar-K-FastICA} & 5.47 & 0.95 & 0.95 & 0.95 & 0.95 \\
MLP & Identity & \textbf{m-ar-K-FastICA} & 1.12 & 0.96 & 0.97 & 0.96 & 0.96 \\
MLP & tanh & \textbf{m-ar-K-FastICA} & 3.40 & 0.97 & 0.97 & 0.97 & 0.97 \\
MLP & ReLU & \textbf{m-ar-K-FastICA} & 2.00 & 0.96 & 0.97 & 0.96 & 0.96
\end{tabular}%
}
\end{table}


\textbf{Table 2.} Results on performance evaluation of baseline (non-optimised) Multi-Layer Perceptron (MLP) in scikit-learn with different kernel functions, to compare the impact of using either the gold standard FastICA feature extraction algorithm or the proposed m-ar-K-FastICA method. The performance of this classifier was assessed on the ‘Heart failure clinical records’ dataset~\citep{chicco2020machine} available in the University California Irvine (UCI) Machine Learning repository.

\begin{table}[H]
\resizebox{\textwidth}{!}{%
\begin{tabular}{llllllll}
\textbf{Classifier} &
  \textbf{Kernel} &
  \textbf{Feature extraction} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Training time \\ \\ (s)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Accuracy \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Precision \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Recall \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}F1-score \\ \\ (0-1)\end{tabular}} \\
MLP & m-arcsinh & FastICA & 0.29 & 0.78 & 0.61 & 0.78 & 0.69 \\
MLP & Identity & FastICA & 0.26 & 0.78 & 0.61 & 0.78 & 0.69 \\
MLP & tanh & FastICA & 0.34 & 0.78 & 0.61 & 0.78 & 0.69 \\
MLP & ReLU & FastICA & 0.30 & 0.78 & 0.61 & 0.78 & 0.69 \\
MLP & m-arcsinh & \textbf{m-ar-K-FastICA} & 0.29 & 0.78 & 0.61 & 0.78 & 0.69 \\
MLP & Identity & \textbf{m-ar-K-FastICA} & 0.47 & 0.78 & 0.61 & 0.78 & 0.69 \\
MLP & tanh & \textbf{m-ar-K-FastICA} & 0.34 & 0.78 & 0.61 & 0.78 & 0.69 \\
MLP & ReLU & \textbf{m-ar-K-FastICA} & 0.38 & 0.78 & 0.61 & 0.78 & 0.69
\end{tabular}%
}
\end{table}

\newpage


\textbf{Table 3.} Results on performance evaluation of baseline (non-optimised) Multi-Layer Perceptron (MLP) in scikit-learn with different kernel functions, to compare the impact of using either the gold standard FastICA feature extraction algorithm or the proposed m-ar-K-FastICA method. The performance of this classifier was assessed on the ‘Haberman survival’ dataset~\citep{Lim:1999} available in the University California Irvine (UCI) Machine Learning repository.

\begin{table}[H]
\resizebox{\textwidth}{!}{%
\begin{tabular}{llllllll}
\textbf{Classifier} &
  \textbf{Kernel} &
  \textbf{Feature extraction} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Training time \\ \\ (s)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Accuracy \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Precision \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Recall \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}F1-score \\ \\ (0-1)\end{tabular}} \\
MLP & m-arcsinh & FastICA & 0.93 & 0.82 & 0.68 & 0.82 & 0.74 \\
MLP & Identity & FastICA & 0.95 & 0.82 & 0.68 & 0.82 & 0.74 \\
MLP & tanh & FastICA & 0.96 & 0.82 & 0.68 & 0.82 & 0.74 \\
MLP & ReLU & FastICA & 1.04 & 0.82 & 0.68 & 0.82 & 0.74 \\
MLP & m-arcsinh & \textbf{m-ar-K-FastICA} & 0.80 & 0.82 & 0.78 & 0.82 & 0.79 \\
MLP & Identity & \textbf{m-ar-K-FastICA} & 1.10 & 0.82 & 0.78 & 0.82 & 0.79 \\
MLP & tanh & \textbf{m-ar-K-FastICA} & 0.92 & 0.82 & 0.78 & 0.82 & 0.79 \\
MLP & ReLU & \textbf{m-ar-K-FastICA} & 1.00 & 0.82 & 0.78 & 0.82 & 0.79
\end{tabular}%
}
\end{table}


\textbf{Table 4.} Results on performance evaluation of baseline (non-optimised) Multi-Layer Perceptron (MLP) in scikit-learn with different kernel functions, to compare the impact of using either the gold standard FastICA feature extraction algorithm or the proposed m-ar-K-FastICA method. The performance of this classifier was assessed on the ‘Parkinson’s’ dataset~\citep{little2007exploiting} available in the University California Irvine (UCI) Machine Learning repository.

\begin{table}[H]
\resizebox{\textwidth}{!}{%
\begin{tabular}{llllllll}
\textbf{Classifier} &
  \textbf{Kernel} &
  \textbf{Feature extraction} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Training time \\ \\ (s)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Accuracy \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Precision \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Recall \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}F1-score \\ \\ (0-1)\end{tabular}} \\
MLP & m-arcsinh & FastICA & 0.80 & 0.77 & 0.59 & 0.77 & 0.67 \\
MLP & Identity & FastICA & 0.78 & 0.77 & 0.59 & 0.77 & 0.67 \\
MLP & tanh & FastICA & 0.80 & 0.77 & 0.59 & 0.77 & 0.67 \\
MLP & ReLU & FastICA & 0.68 & 0.77 & 0.59 & 0.77 & 0.67 \\
MLP & m-arcsinh & \textbf{m-ar-K-FastICA} & 0.67 & 0.77 & 0.59 & 0.77 & 0.67 \\
MLP & Identity & \textbf{m-ar-K-FastICA} & 0.62 & 0.77 & 0.59 & 0.77 & 0.67 \\
MLP & tanh & \textbf{m-ar-K-FastICA} & 0.63 & 0.77 & 0.59 & 0.77 & 0.67 \\
MLP & ReLU & \textbf{m-ar-K-FastICA} & 0.70 & 0.77 & 0.59 & 0.77 & 0.67
\end{tabular}%
}
\end{table}

\newpage


\textbf{Table 5.} Results on performance evaluation of baseline (non-optimised) Multi-Layer Perceptron (MLP) in scikit-learn with different kernel functions, to compare the impact of using either the gold standard FastICA feature extraction algorithm or the proposed m-ar-K-FastICA method. The performance of this classifier was assessed on the ‘SPECTF’ dataset~\citep{Cios:2001} available in the University California Irvine (UCI) Machine Learning repository.

\begin{table}[H]
\resizebox{\textwidth}{!}{%
\begin{tabular}{llllllll}
\textbf{Classifier} &
  \textbf{Kernel} &
  \textbf{Feature extraction} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Training time \\ \\ (s)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Accuracy \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Precision \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}Recall \\ \\ (0-1)\end{tabular}} &
  \textbf{\begin{tabular}[c]{@{}l@{}}F1-score \\ \\ (0-1)\end{tabular}} \\
MLP & m-arcsinh & FastICA & 0.53 & 1.00 & 1.00 & 1.00 & 1.00 \\
MLP & Identity & FastICA & 0.60 & 1.00 & 1.00 & 1.00 & 1.00 \\
MLP & tanh & FastICA & 0.62 & 1.00 & 1.00 & 1.00 & 1.00 \\
MLP & ReLU & FastICA & 0.65 & 1.00 & 1.00 & 1.00 & 1.00 \\
MLP & m-arcsinh & \textbf{m-ar-K-FastICA} & 0.68 & 1.00 & 1.00 & 1.00 & 1.00 \\
MLP & Identity & \textbf{m-ar-K-FastICA} & 0.53 & 1.00 & 1.00 & 1.00 & 1.00 \\
MLP & tanh & \textbf{m-ar-K-FastICA} & 0.65 & 1.00 & 1.00 & 1.00 & 1.00 \\
MLP & ReLU & \textbf{m-ar-K-FastICA} & 0.81 & 1.00 & 1.00 & 1.00 & 1.00
\end{tabular}%
}
\end{table}

\newpage


\vskip 0.3in

\section{Discussion}

\vskip 0.15in

As demonstrated by the highest reliability achieved via the proposed feature extraction method 'm-ar-K-FastICA', quantified in section 2.4, and achieved on 2 datasets assessed (Tables 1 and 3), whilst the same reliability was obtained on the other 3 datasets evaluated, it yielded the best classification performance when considering the same classifier as opposed to the current gold standard FastICA method. For the dataset considered in Table 1, when using the MLP with the identity and ReLU activation functions, leveraging the novel m-ar-K-FastICA method led to 1\% improvement in classification accuracy, recall/sensitivity, and F1-score, whilst the precision increased by 2\%. When considering the dataset in Table 3, precision and F1-score increased by 10\% and 5\% respectively, using the MLP classifier for all activation functions used (m-arcsinh, identity, tanh, ReLU).
\\

Thus, overall, the m-ar-K-FastICA algorithm led to less false positives and increased reliability of the predictions. These improvements in classification performance achieved via the proposed dimensionality reduction method is important to consider for medical applications to reduce psychosocial consequences due to inaccurate screenings~\citep{rasmussen2020psychosocial}, such as those to aid early diagnosis of various tumours, including lung, prostate, and breast cancers.
\\

Besides its higher reliability, leveraging the proposed m-ar-K-FastICA feature extraction method led to reductions in training time for 3 of 5 datasets evaluated, i.e., 14.38\%, 3.27\%, and 1.55\% shorter training times for datasets in Tables 4, 2, and 3 respectively. Thus, also considering the computational efficiency brought by the novel dimensionality reduction method m-ar-K-FastICA, it shows potential to be applied in resource-constrained environments for maximising its positive impact.
\\

Therefore, the m-ar-K-FastICA demonstrates that the m-arcsinh kernel is not only a generalisable function for classification, but also when coupled with feature extraction methods, such as FastICA, to improve the reliability and computational efficiency of classifiers.
As a reliable and computationally efficient approach, the m-ar-K-FastICA is thus deemed a new gold standard dimensionality reduction method, freely available in scikit-learn. 

\newpage


\section{Conclusion}

The m-ar-K-FastICA feature extraction algorithm in scikit-learn provides a new method to achieve dimensionality reduction, even in presence of entropic data. In this study, the proposed approach was shown to lead to higher reliability and computationally efficiency in supervised learning, with regards to MLP-aided classification. As it is made freely available, open source, on the Python and scikit-learn ecosystems, it adds to the choices that both academia and industry can have when selecting a dimensionality reduction method, also in resource-constrained environments. 
\\

Importantly, the proposed m-ar-K-FastICA method was found faster and more reliable for feature extraction than the current gold standard FastICA method. Written in a high-level programming language (Python), it can be used within ML-based pipelines in specific use cases, wherein high precision and reliability need to be obtained, whilst expensive computational hardware may not be available, such as to aid screenings of tumours, including prostate, lung, and breast cancers. Future work involves extending the coupling of the m-arcsinh kernel to benefit other dimensionality reduction methods and provide further choices that the AI community worldwide, including in resource-constrained environments, can have when selecting an extrinsic feature extraction algorithm to aid predictive analytics.


\acks{This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.}


\newpage

\bibliography{ArXiv_paper_LP}

\end{document}